OpenStudy (paki):

19 % students are absent..... 162 are present..... how much are absent ??

OpenStudy (amistre64):

well, if absent + present = 100% .... then present = 100% - absent

OpenStudy (paki):

yeah that is the first step... its answer is 38... how to solve, i dont know...

OpenStudy (amistre64):

100 - 19 is not 38

OpenStudy (anonymous):

81

OpenStudy (amistre64):

yeah .. then 81% of n students = 162

OpenStudy (anonymous):

100-19 = 81

OpenStudy (anonymous):

then find n n=?

OpenStudy (paki):

the question is that if there are 162 students present... and 19% of them are absent... then what is the number of that students....

OpenStudy (amistre64):

...and that is what we are working out :/

OpenStudy (anonymous):

yeah the n

OpenStudy (amistre64):

this might be embarrassing if we all of a sudden started finding tensor values in linear applications.

OpenStudy (amistre64):

not the n, but 19% of the n

OpenStudy (paki):

then

OpenStudy (anonymous):

yeah sorry about that :( 19% of the n

OpenStudy (amistre64):

paki, we need you to participate in the solution process.

OpenStudy (amistre64):

what is the value of n if 81% of n = 162 ?

OpenStudy (anonymous):

@amistre64 already gave you the equation can you solve it now ?

OpenStudy (amistre64):

n/100 = 1% of n so ... 81 * n/100 = 81% of n 81 * n/100 = 162; then solve for 19% of n to determine the number of absents

OpenStudy (anonymous):

how is this going @paki now ?

OpenStudy (paki):

i solved it like this... 162 is equal to 81% and we need 19%... so 162x100/81=200... and 200-162=38... 38 students are absent

OpenStudy (amistre64):

you solved it well then

OpenStudy (amistre64):

200 total .... minus the 162 is fine

OpenStudy (paki):

thanks for helping... :-)

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

so you got it right :) good job @paki

OpenStudy (paki):

thank you guys...

OpenStudy (anonymous):

no problem , good luck :)